what is 7x-5+12^2=3x using the quadratic formula
Answer by pinkgal(1)

7x-5+12^2=3x
To set the left-hand side of the equation equal to 0, move all the expressions to the left-hand side.
7x-5+12^2-3x=0
Since 7x and -3x are like terms, add -3x to 7x to get 4x.
4x-5+12^2=0
Move all terms not containing x to the right-hand side of the equation.
4x=5-12^2
Simplify the right-hand side of the equation.
4x=-139
4x/4=139/4
Simplify the left-hand side of the equation by canceling the common factors.
x=-139/4